OpenStudy (anonymous):

integral 1/(1-2x)^(3/2)= -2/sqrt(1-2x) true or false

OpenStudy (anonymous):

\[\int\limits dx/\sqrt[3]{(1-2x)^2}= -2/\sqrt{(1-2x)} TRUEE??\]

OpenStudy (bahrom7893):

wait ur original question says ^3/2 the one below says ^2/3

OpenStudy (anonymous):

hahah sorry focus on the original

OpenStudy (bahrom7893):

okay..

OpenStudy (bahrom7893):

false

OpenStudy (bahrom7893):

Please click on Good Answer if I helped, I really want to get to the next level, thanks =)!